# Euclid Quick Data Release (Q1)

### v1

@@ 4.1 Photometric properties
The visually identified quasar sample contains 3468 sources (38 % of the entire sample) covering the redshift range of $0<z\lesssim 4.8$, which are compiled into a catalogue detailed in Table 5. This catalogue includes source IDs and coordinates, redshifts, spectroscopic quality indicators, PSF fraction measurements from VIS imaging (Q1-SP015) for a redshift-limited subsample ($0.5<z<2$), and magnitudes across the Euclid bands. Among them, 2753 are Gaia sources and 715 are non-Gaia sources.
The redshift and magnitude distributions of the full sample and Gaia/non-Gaia subsets are shown in Fig. 3. The redshift density distributions peak in the range $0.89\lesssim z\lesssim 1.83$ when the H $\alpha$ emission line is in the observed wavelength range, and drop sharply at $z\approx 0.89$ and $z\approx 1.83$ when H $\alpha$ moves out of the wavelength range. As shown in the histograms of Euclid magnitudes, this visually identified quasars represent a bright sample with a median $I_{\scriptscriptstyle\rm E}$ of 20.5, and median magnitudes of 19.9 in $Y_{\scriptscriptstyle\rm E}$, 19.7 in $J_{\scriptscriptstyle\rm E}$, and 19.5 in $H_{\scriptscriptstyle\rm E}$. The non-Gaia subset is overall 2 magnitudes fainter than the Gaia subset in $I_{\scriptscriptstyle\rm E}$, 0.8 magnitudes fainter in $J_{\scriptscriptstyle\rm E}$ and 0.6 magnitudes fainter in $H_{\scriptscriptstyle\rm E}$. The faintest identified quasars in this work have $I_{\scriptscriptstyle\rm E}\approx 27$, $Y_{\scriptscriptstyle\rm E}\approx 23$, and $J_{\scriptscriptstyle\rm E}\approx H_{\scriptscriptstyle\rm E}\approx 22.5$.
@@
Figure 4: Left: Distribution of the median spectral S/N within [12 047, 18 734] Å for visually identified quasars. The red dashed line at $\mathrm{S/N}=2$ marks an empirical threshold below which the number of successful identifications declines rapidly. Middle and Right: Median S/N versus $J_{\scriptscriptstyle\rm E}$ and $H_{\scriptscriptstyle\rm E}$ magnitudes, respectively. Blue points show individual quasars, and the black lines indicate linear fits to $\log_{10}(\mathrm{S/N})$, with the fitted relations annotated. The red dashed line again marks $\mathrm{S/N}=2$, and its intersection with the fit defines empirical limiting magnitudes of $J_{\scriptscriptstyle\rm E}\approx 21.5$ and $H_{\scriptscriptstyle\rm E}\approx 21.3$, shown by green dashed lines. These values represent practical limits for reliable spectral identification of quasars in Q1.
@@
These point-like criteria are similar to the definition of Euclid point-like sources in Q1-SP027.
The Gaia and non-Gaia quasar subsets share nearly identical colour spaces on the W1W2W1-W2 versus W2W3W2-W3 diagram (Fig. 5), which have been described as AGN regions by many previous studies (e.g., 2012ApJ...753...30S; 2012AJ....144...49W; 2012MNRAS.426.3271M; 2018ApJS..234...23A).
On Euclid colour planes (Fig. 6), however, the Gaia and non-Gaia quasar samples become more separable. The non-Gaia subset occupies colour spaces redder than the Gaia subset, while partly overlapping. The optical-faint and infrared-bright selection of the non-Gaia subset makes it a representative sample of red quasars. As shown in Fig. 6, the Euclid colour cuts for red quasars ($Y_{\scriptscriptstyle\rm E}-H_{\scriptscriptstyle\rm E}>0.7$, $J_{\scriptscriptstyle\rm E}-H_{\scriptscriptstyle\rm E}>0.3$, and $I_{\scriptscriptstyle\rm E}-H_{\scriptscriptstyle\rm E}>1.8$) from Q1-SP023 select the redder half of non-Gaia quasars, which only have a small overlap with the Gaia quasars.
@@
Figure 6: Euclid and external (Hyper Suprime-Cam, HSC; 2018PASJ...70S...1M; 2018PASJ...70S...4A) colour-colour diagrams of spectroscopically identified quasars in this work, where Gaia-detected sources are shown as blue open squares, and sources not in Gaia DR3 are shown as orange circles. The density distribution of the Gaia-detected sample is indicated with blue contour lines, and that of the non-Gaia sample is indicated with red contour lines. The density of Euclid point-like sources is shown as grey hexagonal binning plots in the background.
@@ 4.2 A stacked emission-line map and the redshift challenge
To ensure robust composite building and follow-up spectral analysis, we select a golden sample of 2868 visually identified quasars with the following constraints: (i) median S/N of the spectrum higher than 3 within [12 047, 18 734] Å; and (ii) containing $\leq 15$ invalid pixels (NaN, or zero flux values).
To better understand the systematic effects of the redshift determination using ${\rm RG}_{\scriptscriptstyle\rm E}$ spectra, we construct a stacked emission-line map of the golden sample on the redshift–wavelength plane (Fig. 7) after normalising each spectrum with a percentile-based scaling:
@@
From Fig. 7, the most prominent emission line among all spectra is H $\alpha$, which lies in the wavelength range of [12 047, 18 734] Å at redshift $0.83\lesssim z\lesssim 1.85$. Other prominent emission lines include He i + Pa $\gamma$ at $0.11<z<0.73$, H $\beta$ + [O iii] at $1.5\lesssim z\lesssim 2.8$, and Mg ii at $z>3.3$. While present in the emission-line map, H $\gamma$ becomes faint at high redshift, especially when H $\beta$ + [O iii] move out of the observed wavelength range. The low equivalent width (EW) and S/N of H $\gamma$ pose a challenge to the redshift determination at $2.8<z<3.3$, which can only be alleviated by combining with data at other wavelengths. In future data releases of Euclid, the blue-grism data of NISP, covering 926–1366 nm, will be available for Euclid Deep and Auxiliary fields (EuclidSkyOverview). Inclusion of the blue-grism data will increase the efficiency of source identification and redshift determination in these deep fields.
@@ 4.3 Composite Euclid spectra of the golden sample of bright quasars
A mean or median composite spectrum of a sample of quasars is useful for understanding the average spectral properties across a wide wavelength range, and for providing a representative spectrum template (e.g., 2001AJ....122..549V; Glikman_2006).
To generate the mean and median composite quasar spectra spanning the rest-frame optical to NIR wavelengths, we adopted a procedure similar to that described by 2001AJ....122..549V, with a 1D ‘drizzle’ technique to improve the sampling of the low-resolution spectra. We briefly introduce the procedure below, and refer to Appendix B for technical details. We first correct the Milky Way dust extinction of the spectra using E(BV)E(B-V) values from the Galactic dust map produced by 2016A&A...596A.109P, and the extinction law of 2023ApJ...950...86G assuming $R_{V}=3.1$. The packages dustmaps (2018JOSS....3..695G) and dust_extinction (2024JOSS....9.7023G) are used for the correction. The spectra are then sorted in ascending order of redshift. The spectrum with the lowest redshift is normalised arbitrarily to have a unit mean flux density. Each subsequent spectrum is shifted to the rest-frame based on its measured redshift and normalised to match the mean flux density of the mean composite built from all lower-redshift spectra, within their overlapping wavelength range.
All spectra are resampled onto a common rest-frame wavelength grid with a constant bin size of $\Delta\lambda=4$ Å, using a flux-conserving re-binning method to preserve the integrated flux density in each bin. This process is mathematically equivalent to a 1D version of the Drizzle algorithm (2002PASP..114..144F), which has been widely applied to image reconstructions by combining dithered, undersampled images in HST and JWST surveys (e.g., 2011ApJS..197...36K; 2023ApJS..268...64W; 2023ApJ...946L..12B). We set $\Delta\lambda=4$ Å to match the smallest native rest-frame pixel size near the blue end while maintaining higher per-pixel S/N at the highest redshifts. For reference, the observed sampling of 13.4 Å at $z=2.35$ corresponds to $13.4\,\AA /(1+2.35)=4$ Å in the rest frame, and the blue-end wavelength at this redshift is $12047\,\AA /(1+2.35)=3596$ Å. Consequently, the common 4-Å grid oversamples the data at longer rest wavelengths and applies mild downsampling at the shortest rest wavelengths contributed by the highest-redshift quasars. By stacking the oversampled rest-frame spectra of different redshifts through drizzle, the composite recovers some information that is lost in the individual spectra at rest-frame $\lambda>3596\,\AA$ , giving finer emission line features than individual ones.
@@
Table 2: Mean, geometric mean, and median Euclid Q1 quasar composite spectra, along with RMS, S/N and number of spectra in each wavelength bin ($N_{\mathrm{spec}}$). [^5] Table 2 is available in its entirety in the electronic edition. A portion is shown here for guidance regarding its form and content.
@@
240.0 $57\pm 0.90 13.13$$13\pm 0.97$ 14.26 3.49 15.09 7
240.4 $13.36\pm 0.86$ $12.90\pm 0.83$ 13.47 3.23 15.48 7
240.8 $14.32\pm 0.98$ $13.81\pm 1.01$ 13.21 3.69 14.55 7
241.2 $14.47\pm 0.95$ $13.96\pm 0.99$ 14.36 3.97 15.30 7
241.6 $13.02\pm 0.90$ $12.36\pm 1.16$ 14.36 3.55 14.43 7
1698.4 $0.83\pm 0.02$ $0.83\pm 0.02$ 0.84 0.06 55.18 9
1698.8 $0.83\pm 0.02$ $0.82\pm 0.02$ 0.83 0.06 50.32 9
1699.2 $0.77\pm 0.02$ $0.72\pm 0.02$ 0.83 0.20 46.34 9
1699.6 $0.84\pm 0.02$ $0.84\pm 0.02$ 0.83 0.05 48.50 8
1700.0 $0.86\pm 0.02$ $0.86\pm 0.02$ 0.86 0.08 50.61 8
Figure 8 presents the mean and median composite spectra derived from our sample, together with the RMS scatter around the mean composite. The mean quasar composite spectrum from Glikman_2006, the mean composite of type 1 AGN from EP-Lusso, and a mean composite spectrum constructed by EP-Lusso using data sets from 2008ApJS..174..282L; 2011MNRAS.414..218L; 2013MNRAS.432..113L, are also shown for comparison. Prominent emission lines are marked in the composite spectra for reference, including Mg ii $\lambda 2800$, [O ii] $\lambda 3728$, H $\gamma$, H $\beta$, [O iii] $\lambda\lambda 4960,5008$, H $\alpha$, He i, Pa $\gamma$, and Pa $\beta$. The mean and median composite spectra are consistent with each other.
@@
Figure 8: Mean and median composite quasar spectra constructed from our golden sample (black and red curves, respectively), with the RMS scatter around the mean composite indicated by the blue-shaded region. The quasar composite spectrum from Glikman_2006 is plotted in magenta, the type 1 AGN composite from EP-Lusso is plotted in green, and the mean composite constructed using data from 2008ApJS..174..282L; 2011MNRAS.414..218L; 2013MNRAS.432..113L is plotted in orange. Prominent emission lines are marked for reference.
@@ 5.1 Morphological properties of the bright quasar sample
With a pixel scale of 0.1 and stable PSF of VIS imaging, Euclid resolves obvious galaxy structure from the local Universe to at least $z=1.5$ (with particularly rich statistics at $0.3<z<0.7$; see e.g. Q1-SP047). These data enable the detailed structural characterisation of quasar hosts through uniform measurements of Srsic-based and model-independent parameters (Q1-SP040; Q1-TP004), as well as the separation of compact nuclear light from extended galaxy emission (Q1-SP015).
@@
To take advantage of the detailed structural information provided by the VIS images, we examine two redshift regimes separately: (i) a low-redshift ($z<0.5$) subset with 341 sources, and (ii) an intermediate-redshift ($0.5<z<2$) subset with 2361 sources. Both subsets are selected to have valid morphological parameters measured from VIS imaging, including parameters from the single-component Srsic fit (Q1-SP040): Srsic index ($n_{{\scriptscriptstyle\rm VIS}}$), effective radius ($R_{\mathrm{e,}{\scriptscriptstyle\rm VIS}}$), and axis ratio ($\rho_{{\scriptscriptstyle\rm VIS}}$); and the model-independent CAS parameters (2003ApJS..147....1C; Q1-TP004): concentration, asymmetry, and clumpiness (smoothness). The intermediate-redshift subset is further supplemented with a VIS PSF fraction ($f_{{\scriptscriptstyle\rm PSF}}$), derived by the deep-learning-based $f_{{\scriptscriptstyle\rm PSF}}$ prediction model trained on simulated galaxy images with different levels of $f_{{\scriptscriptstyle\rm PSF}}$ added to them (Q1-SP015). We choose the redshift range $0.5<z<2$ because the model was trained and validated in this range.
@@
As shown in Fig. 11 and Table 4, the low-redshift population shows a large source extent with median $\mu_{\mathrm{max}}-\mathrm{mag}\ =-1.67$, significantly higher than the median $\mu_{\mathrm{max}}-\mathrm{mag}\ =-2.87$ of the intermediate-redshift sample. The $\mu_{\mathrm{max}}-\mathrm{mag}$ value anticorrelates with the Srsic index and concentration. In particular, the most compact sources in this population ($\mu_{\mathrm{max}}-\mathrm{mag}\approx-2.5$) have a Srsic index around 5.5, which is the upper limit set by Q1-SP040. They have suggested that fits with Srsic indices above 5.45 should be removed from any Srsic-based analysis. In this low-redshift population, 50 % of the sources have $n_{{\scriptscriptstyle\rm VIS}}$ higher than 5.45. Such near-boundary Srsic indices indicate that the single-component Srsic model cannot describe the light profiles of AGN with bright cores. When requiring $n_{{\scriptscriptstyle\rm VIS}}<5.45$, the median $n_{{\scriptscriptstyle\rm VIS}}$ is 2.31, which is still significantly higher than the peak value of 0.8 among all galaxies in Q1-SP040. The median $R_{\mathrm{e,}{\scriptscriptstyle\rm VIS}}$ of $n_{{\scriptscriptstyle\rm VIS}}<5.45$ sources is 1.0, identical to the peak value of $R_{\mathrm{e,}{\scriptscriptstyle\rm VIS}}$ in Q1-SP040.
@@
The CAS parameters further characterise the resolved structure of the low-redshift quasars. The median concentration ($C=4.27$) is high compared to the general galaxy population (e.g., $C\approx 2.5$, see Q1-SP040) but with significant scatter, reflecting the coexistence of bright nuclear light and extended hosts. Asymmetry values are moderate (median $A=0.36$) with a tail to $A>1$, indicating disturbed morphologies or nearby companions. The clumpiness parameter is elevated (median $S=0.19$), suggesting clumpy emission in many cases, possibly due to ongoing star formation (e.g. 2003ApJS..147....1C; 2014ARA&A..52..291C). Overall, at $z<0.5$, the host galaxies of bright quasars are frequently resolved and often display interacting features and star-forming clumps.
Table 4: Median VIS morphology parameters and $f_{{\scriptscriptstyle\rm PSF}}$. $f_{{\scriptscriptstyle\rm PSF}}$ is available only for $0.5<z<2$.
@@
Figure 11: Corner plot showing the joint distributions of morphological parameters measured from Euclid VIS imaging for 341 sources at $z<0.5$. Displayed parameters include visual-inspection redshift ($z_{\mathrm{vi}}$), Srsic index ($n_{{\scriptscriptstyle\rm VIS}}$), axis ratio ($\rho_{{\scriptscriptstyle\rm VIS}}$), logarithmic effective radius ($\log_{10}(R_{\mathrm{e,}{\scriptscriptstyle\rm VIS}}/1^{\prime\prime})$), concentration, asymmetry, clumpiness, and $\mu_{\mathrm{max}}-\mathrm{mag}$.

### v2

@@ 4.1 Photometric properties
The 3468 sources of the visually identified quasar sample are compiled into a catalogue detailed in Table 5. This catalogue includes source IDs and coordinates, redshifts, spectroscopic quality indicators, point spread function (PSF) fraction measurements from VIS imaging (Euclid Collaboration: Margalef-Bentabol et al. 2025) for a redshift-limited subsample ($0.5<z<2$), and magnitudes across the Euclid bands. Among them, 2753 are Gaia sources and 715 are non-Gaia sources.
The redshift and magnitude distributions of the full sample, and Gaia or non-Gaia subsets, are shown in Fig. 3. The redshift density distributions peak in the range $0.89\lesssim z\lesssim 1.83$ when the H $\alpha$ emission line is in the observed wavelength range, and drop sharply at $z\approx 0.89$ and $z\approx 1.83$ when H $\alpha$ moves out of the wavelength range. As shown in the histograms of Euclid magnitudes, these visually identified quasars represent a bright sample with a median $I_{\scriptscriptstyle\rm E}$ of 20.5, and median magnitudes of 19.9 in $Y_{\scriptscriptstyle\rm E}$, 19.7 in $J_{\scriptscriptstyle\rm E}$, and 19.5 in $H_{\scriptscriptstyle\rm E}$. The non-Gaia subset is overall 2 magnitudes fainter than the Gaia subset in $I_{\scriptscriptstyle\rm E}$, 0.8 magnitudes fainter in $J_{\scriptscriptstyle\rm E}$ and 0.6 magnitudes fainter in $H_{\scriptscriptstyle\rm E}$. The faintest identified quasars in this work have $I_{\scriptscriptstyle\rm E}\approx 27$, $Y_{\scriptscriptstyle\rm E}\approx 23$, and $J_{\scriptscriptstyle\rm E}\approx H_{\scriptscriptstyle\rm E}\approx 22.5$.
@@
Figure 4: Left: Distribution of the median spectral S/N within [12 047, 18 734] Å for visually identified quasars. The dashed red line at $\mathrm{S/N}=2$ marks an empirical threshold below which the number of successful identifications declines rapidly. Middle and Right: Median S/N versus $J_{\scriptscriptstyle\rm E}$ and $H_{\scriptscriptstyle\rm E}$ magnitudes, respectively. Blue points show individual quasars, and the black lines indicate linear fits to $\log_{10}(\mathrm{S/N})$, with the fitted relations annotated. The dashed red line again marks $\mathrm{S/N}=2$, and its intersection with the fit defines empirical limiting magnitudes of $J_{\scriptscriptstyle\rm E}\approx 21.5$ and $H_{\scriptscriptstyle\rm E}\approx 21.3$, shown by dashed green lines. These values represent practical limits for reliable spectral identification of quasars in Q1.
@@
These point-like criteria are similar to the definition of Euclid point-like sources in Euclid Collaboration: Matamoro Zatarain et al. (2025).
The Gaia and non-Gaia quasar subsets share nearly identical colour spaces on the W1W2W1-W2 versus W2W3W2-W3 diagram (Fig. 5). These have been described as AGN regions by many previous studies (e.g. Stern et al. 2012; Wu et al. 2012; Mateos et al. 2012; Assef et al. 2018). On Euclid colour planes (Fig. 6), however, the Gaia and non-Gaia quasar samples become more separable. The non-Gaia subset occupies colour spaces redder than the Gaia subset, while partly overlapping. The optical-faint and infrared-bright selection of the non-Gaia subset makes it a representative sample of red quasars. As shown in Fig. 6, the Euclid colour cuts for red quasars ($Y_{\scriptscriptstyle\rm E}-H_{\scriptscriptstyle\rm E}>0.7$, $J_{\scriptscriptstyle\rm E}-H_{\scriptscriptstyle\rm E}>0.3$, and $I_{\scriptscriptstyle\rm E}-H_{\scriptscriptstyle\rm E}>1.8$) from Euclid Collaboration: Tarsitano et al. (2025, hereafter T25) select the redder half of non-Gaia quasars, which only have a small overlap with the Gaia quasars.
@@
Figure 6: Euclid and external (Hyper Suprime-Cam, HSC; Miyazaki et al. 2018; Aihara et al. 2018) colour-colour diagrams of spectroscopically identified quasars in this work, where Gaia-detected sources are shown as blue open squares, and sources not in Gaia DR3 are shown as orange circles. The density distribution of the Gaia-detected sample is indicated with blue contour lines, and that of the non-Gaia sample is indicated with red contour lines. The density of Euclid point-like sources is shown as grey hexagonal binning plots in the background.
@@ 4.2 A stacked emission-line map and the redshift challenge
To ensure robust composite building and follow-up spectral analysis, we selected a golden sample of 2868 visually identified quasars with the following constraints: (i) median S/N of the spectrum higher than 3 within [12 047, 18 734] Å; and (ii) containing $\leq 15$ invalid pixels (NaN, or zero flux values).
To better understand the systematic effects of the redshift determination using ${\rm RG}_{\scriptscriptstyle\rm E}$ spectra, we constructed a stacked emission-line map of the golden sample on the redshift–wavelength plane (Fig. 7) after normalising each spectrum with a percentile-based scaling,
@@
From Fig. 7, the most prominent emission line among all spectra is H $\alpha$, which lies in the wavelength range of [12 047, 18 734] Å at redshift $0.83\lesssim z\lesssim 1.85$. Other prominent emission lines include He i + Pa $\gamma$ at $0.11<z<0.73$, H $\beta$ + [O iii] at $1.5\lesssim z\lesssim 2.8$, and Mg ii at $z>3.3$. While present in the emission-line map, H $\gamma$ becomes faint at high redshift, especially when H $\beta$ + [O iii] move out of the observed wavelength range. The low equivalent width (EW) and S/N of H $\gamma$ pose a challenge to the redshift determination at $2.8<z<3.3$, which can only be alleviated by combining with data at other wavelengths. In future data releases of Euclid, the blue-grism data of NISP, covering 926–1366 nm, will be available for Euclid Deep and Auxiliary fields (Euclid Collaboration: Mellier et al. 2025). Inclusion of the blue-grism data will increase the efficiency of source identification and redshift determination in these deep fields.
@@ 4.3 Composite Euclid spectra of the golden sample of bright quasars
A mean or median composite spectrum of a sample of quasars is useful for understanding the average spectral properties across a wide wavelength range, and for providing a representative spectrum template (e.g. Vanden Berk et al. 2001; Glikman et al. 2006). To generate the mean and median composite quasar spectra spanning the rest-frame optical to NIR wavelengths, we adopted a procedure similar to that described by Vanden Berk et al. (2001), with a 1D ‘drizzle’ technique to improve the sampling of the low-resolution spectra. We briefly introduce the procedure below, and refer to Appendix C for technical details. We first corrected the Milky Way dust extinction of the spectra using E(BV)E(B-V) values from the Galactic dust map produced by Planck Collaboration et al. (2016), and the extinction law of Gordon et al. (2023) assuming $R_{V}=3.1$. The packages dustmaps (Green 2018) and dust_extinction (Gordon 2024) are used for the correction. The spectra are then sorted in ascending order of redshift. The spectrum with the lowest redshift is normalised arbitrarily to have a unit mean flux density. Each subsequent spectrum is shifted to the rest-frame based on its measured redshift and normalised to match the mean flux density of the mean composite built from all lower-redshift spectra, within their overlapping wavelength range.
All spectra were resampled onto a common rest-frame wavelength grid with a constant bin size of $\Delta\lambda=4$ Å, using a flux-conserving re-binning method to preserve the integrated flux density in each bin. This process is mathematically equivalent to a 1D version of the Drizzle algorithm (Fruchter and Hook 2002), which has been widely applied to image reconstructions by combining dithered, undersampled images in HST and JWST surveys (e.g. Koekemoer et al. 2011; Williams et al. 2023; Bagley et al. 2023). We set $\Delta\lambda=4$ Å to match the smallest native rest-frame pixel size near the blue end while maintaining higher per-pixel S/N at the highest redshifts. For reference, the observed sampling of 13.4 Å at $z=2.35$ corresponds to $13.4\,\AA /(1+2.35)=4$ Å in the rest frame, and the blue-end wavelength at this redshift is $12047\,\AA /(1+2.35)=3596$ Å. Consequently, the common 4-Å grid oversamples the data at longer rest wavelengths and applies mild downsampling at the shortest rest wavelengths contributed by the highest-redshift quasars. By stacking the oversampled rest-frame spectra of different redshifts through drizzle, the composite recovers some information that is lost in the individual spectra at rest-frame $\lambda>3596\,\AA$ , giving finer emission line features than individual ones.
@@
Table 2: Mean, geometric mean, and median Euclid Q1 quasar composite spectra, along with RMS, S/N and the number of spectra in each wavelength bin ($N_{\mathrm{spec}}$). [^5] Table 2 is available in its entirety at the CDS. A portion is shown here for guidance regarding its form and content.
@@
240.0 $15.00\pm 1.08$ $14.71\pm 1.13$ 14.35 2.97 13.83 5
240.4 $13.11\pm 1.03$ $12.51\pm 0.96$ 13.44 3.70 12.74 5
240.8 $15.90\pm 1.22$ $15.65\pm 1.18$ 16.19 2.83 12.99 5
241.2 $14.82\pm 1.16$ $14.13\pm 1.26$ 14.39 4.59 12.73 5
241.6 $14.93\pm 1.06$ $14.87\pm 1.05$ 14.59 1.35 14.12 5
1698.4 $0.83\pm 0.01$ $0.83\pm 0.02$ 0.84 0.06 55.20 9
1698.8 $0.83\pm 0.02$ $0.82\pm 0.02$ 0.83 0.06 50.31 9
1699.2 $0.77\pm 0.02$ $0.72\pm 0.02$ 0.83 0.20 46.35 9
1699.6 $0.84\pm 0.02$ $0.84\pm 0.02$ 0.83 0.05 48.51 8
1700.0 $0.86\pm 0.02$ $0.86\pm 0.02$ 0.86 0.08 50.59 8
Figure 8 presents the mean and median composite spectra derived from our sample, together with the RMS scatter around the mean composite. The mean quasar composite spectrum from Glikman et al. (2006), the mean composite of type 1 AGNs from Euclid Collaboration: Lusso et al. (2024), and a mean composite spectrum constructed by Euclid Collaboration: Lusso et al. (2024) using datasets from Landt et al. (2008, 2011, 2013), are also shown for comparison. Prominent emission lines are marked in the composite spectra for reference, including Mg ii $\lambda 2800$, [O ii] $\lambda 3728$, H $\gamma$, H $\beta$, [O iii] $\lambda\lambda 4960,5008$, H $\alpha$, He i, Pa $\gamma$, and Pa $\beta$. The mean and median composite spectra are consistent with each other.
@@
Figure 8: Mean and median composite quasar spectra constructed from our golden sample (black and red curves, respectively), with the RMS scatter around the mean composite indicated by the shaded blue region. The quasar composite spectrum from Glikman et al. (2006) is plotted in magenta, the type 1 AGN composite from Euclid Collaboration: Lusso et al. (2024) is plotted in green, and the mean composite constructed using data from Landt et al. (2008, 2011, 2013) is plotted in orange. Prominent emission lines are marked for reference.
@@ 5.1 Morphological properties of the bright quasar sample
With a pixel scale of 0.1 and stable PSF of VIS imaging, Euclid resolves obvious galaxy structure from the local Universe to at least $z=1.5$ (with particularly rich statistics at $0.3<z<0.7$; see e.g. Euclid Collaboration: Walmsley et al. 2025). These data enable the detailed structural characterisation of quasar hosts through uniform measurements of Sérsic-based and model-independent parameters (Euclid Collaboration: Quilley et al. 2025; Euclid Collaboration: Romelli et al. 2025), as well as the separation of compact nuclear light from extended galaxy emission (Euclid Collaboration: Margalef-Bentabol et al. 2025).
@@
To take advantage of the detailed structural information provided by the VIS images, we examined two redshift regimes separately: (i) a low-redshift ($z<0.5$) subset with 341 sources, and (ii) an intermediate-redshift ($0.5<z<2$) subset with 2361 sources. Both subsets are selected to have valid morphological parameters measured from VIS imaging, including parameters from the single-component Sérsic fit (Euclid Collaboration: Quilley et al. 2025): Sérsic index ($n_{{\scriptscriptstyle\rm VIS}}$), effective radius ($R_{\mathrm{e,}{\scriptscriptstyle\rm VIS}}$), and axis ratio ($\rho_{{\scriptscriptstyle\rm VIS}}$); and the model-independent CAS parameters (Conselice 2003; Euclid Collaboration: Romelli et al. 2025): concentration, asymmetry, and clumpiness (smoothness). The intermediate-redshift subset is further supplemented with a VIS PSF fraction ($f_{{\scriptscriptstyle\rm PSF}}$), derived by the deep-learning-based $f_{{\scriptscriptstyle\rm PSF}}$ prediction model trained on simulated galaxy images with different levels of $f_{{\scriptscriptstyle\rm PSF}}$ added to them (Euclid Collaboration: Margalef-Bentabol et al. 2025). We choose the redshift range $0.5<z<2$ because the model was trained and validated in this range.
@@
As shown in Fig. 11 and Table 4, the low-redshift population shows a large source extent with median $\mu_{\mathrm{max}}-\mathrm{mag}\ =-1.67$, significantly higher than the median $\mu_{\mathrm{max}}-\mathrm{mag}\ =-2.87$ of the intermediate-redshift sample. The $\mu_{\mathrm{max}}-\mathrm{mag}$ value anticorrelates with the Sérsic index and concentration. In particular, the most compact sources in this population ($\mu_{\mathrm{max}}-\mathrm{mag}\approx-2.5$) have a Sérsic index around 5.5, which is the upper limit set by Euclid Collaboration: Quilley et al. (2025). They have suggested that fits with Sérsic indices above 5.45 should be removed from any Sérsic-based analysis. In this low-redshift population, 50 % of the sources have $n_{{\scriptscriptstyle\rm VIS}}$ higher than 5.45. Such near-boundary Sérsic indices indicate that the single-component Sérsic model cannot describe the light profiles of AGNs with bright cores. When requiring $n_{{\scriptscriptstyle\rm VIS}}<5.45$, the median $n_{{\scriptscriptstyle\rm VIS}}$ is 2.31, which is still significantly higher than the peak value of 0.8 among all galaxies in Euclid Collaboration: Quilley et al. (2025). The median $R_{\mathrm{e,}{\scriptscriptstyle\rm VIS}}$ of $n_{{\scriptscriptstyle\rm VIS}}<5.45$ sources is 1.0, identical to the peak value of $R_{\mathrm{e,}{\scriptscriptstyle\rm VIS}}$ in Euclid Collaboration: Quilley et al. (2025).
@@
The CAS parameters further characterise the resolved structure of the low-redshift quasars. The median concentration ($C=4.27$) is high compared to the general galaxy population (e.g. $C\approx 2.5$, see Euclid Collaboration: Quilley et al. 2025) but with significant scatter, reflecting the coexistence of bright nuclear light and extended hosts. Asymmetry values are moderate (median $A=0.36$) with a tail to $A>1$, indicating disturbed morphologies or nearby companions. The clumpiness parameter is elevated (median $S=0.19$), suggesting clumpy emission in many cases, possibly due to ongoing star formation (e.g. Conselice 2003, 2014). Overall, at $z<0.5$, the host galaxies of bright quasars are frequently resolved and often display interacting features and star-forming clumps.
[…]
@@
Figure 11: Corner plot showing the joint distributions of morphological parameters measured from Euclid VIS imaging for 341 sources at $z<0.5$. Displayed parameters include visual-inspection redshift ($z_{\mathrm{vi}}$), Sérsic index ($n_{{\scriptscriptstyle\rm VIS}}$), axis ratio ($\rho_{{\scriptscriptstyle\rm VIS}}$), logarithmic effective radius ($\log_{10}(R_{\mathrm{e,}{\scriptscriptstyle\rm VIS}}/1^{\prime\prime})$), concentration ($C$), asymmetry (AA), clumpiness ($S$), and $\mu_{\mathrm{max}}-\mathrm{mag}$.